Suppose a population consists of the five measurements: 2, 6, 8, 0, and 1:



1. What is the mean and standard deviation of the population?



2. How many different samples of size n=2 can be drawn from the population? List them with their corresponding means.



Follow the step by step procedure.

1
Expert's answer
2022-04-22T02:39:34-0400

"\mu =(2+6+8+0+1)/5=3.4"

"\sigma=\sqrt \frac{\sum(x-\mu)^2}{n}=\sqrt \frac{1.96+6.76+21.16+11.56+5.76}{5}=3.1"

2. "C^2_5=\frac{5!}{3!2!}=10" Number of samples

m(2,6)=(2+6)/2=4

m(2,8)=(2+8)/2=5

m(2,0)=(2+0)/2=1

m(2,1)=(2+1)/2=1.5

m(6,8)=(6+8)/2=7

m(6,0)=(6+0)/2=3

m(6,1)=(6+1)/2=3.5

m(8,0)=(8+0)/2=4

m(8,1)=(8+1)/2=4.5

m(0,1)=(0+1)/2=0.5